# Interpretable Time Series Autoregression for Periodicity Quantification

### v1

@@ Abstract
Time series autoregression is a classical statistical model for capturing auto-correlations and identifying temporal patterns such as periodicity and seasonality. In this work, we propose a novel sparse autoregression framework from an interpretable machine learning perspective and the model interpretability for periodicity quantification is reinforced by $\ell_{0}$-norm induced sparsity constraints. On the time-varying time series data, we reformulate the sparse autoregression and convert the involved optimization problem into a mixed-integer optimization (MIO). To accelerate it, we develop a subspace pursuit based decision variable pruning (DVP) strategy to reduce the search space. On the multidimensional time series that involves complicated spatial and temporal dimensions, we propose a spatially- and time-varying sparse autoregression model and resolve the corresponding MIO problem by developing a two-stage optimization scheme. In particular, the proposed scheme makes the model scalable to large problems even with millions of decision variables. Empirically, we conduct extensive experiments to evaluate the proposed models on real-world time series data. First, we demonstrate that the MIO solver can be drastically accelerated through the DVP strategy, while maintaining the same solution quality as a full MIO solver. Applying the time-varying sparse autoregression model to ridesharing trip data, we uncover both daily and weekly periodicities and reveal long-term changes in regularity of human mobility. Second, we demonstrate the spatial patterns of yearly seasonality in climate variable time series such as temperature and precipitation across the past four decades, and our model allows to discover dynamic climate patterns and identify climate phenomena such as El Niño in sea surface temperature. Thus, our framework lays an insightful foundation for understanding periodicity and seasonality in real-world time series.
@@ 1 Introduction
Many real-world systems exhibit complex temporal patterns, including periodicity, seasonality, and anomalies. Detecting and understanding these patterns is essential for anticipating system behavior, identifying disruptions, and supporting operational decision-making. In dynamic environments such as urban transportation and climate systems, periodicities can shift due to external factors—policy interventions, demand changes, environmental variability, global events, or extreme climate phenomena—making interpretable, adaptive, and data-driven methods indispensable. A central challenge is to automatically identify dominant periodic components from spatially- and time-varying systems, track their evolution over a long-term time period, and distinguish true structural changes from random variability.
Urban transportation systems display strong periodicity driven by commuting patterns, business cycles, and travel demand. Fig. 1 shows the regularity of ridesharing activity in Chicago, highlighting its weekly mobility patterns. However, such patterns are not fixed—they evolve with infrastructure changes, economic conditions, and disruptive events. For example, the COVID-19 pandemic in 2020 led to a collapse in established mobility periodicities due to lockdowns and shifts to remote work. These disruptions raise key questions: How do periodic structures evolve over time? Can we systematically quantify such changes in a time-varying system? Addressing these questions is critical for planning, forecasting, and adaptive resource allocation.
@@
To address these needs, we develop a unified interpretable machine learning framework for identifying and quantifying time-varying periodicity and structural shifts in real-world time series. Our models build on classical autoregression (AR) [4, 5] but incorporate sparse structure via $\ell_{0}$-norm induced sparsity constraints to promote interpretability. Inspired by recent advances in sparse regression [6] and time series convolution [7], we formulate interpretable AR models capable of isolating dominant periodic patterns over time and space. To the best of our knowledge, this represents the first application of exact sparse AR to the task of periodicity quantification in real-world complex time series. Overall, this work makes the following contributions:
@@
  • Time-Varying SAR (TV-SAR): We extend SAR to characterize non-stationary time series, enforcing consistent support sets across time segments to enhance interpretability. To improve scalability, we introduce a decision variable pruning (DVP) strategy that narrows the MIO search space using fast greedy approximations such as subspace pursuit.
  • Spatially- and Time-Varying SAR (STV-SAR): We propose a scalable model for multidimensional time series that vary over both space and time. A two-stage optimization procedure—global support set selection via MIO, followed by local coefficient estimation via quadratic optimization—makes the method tractable for millions of decision variables.
@@ 2 Literature Review
2.1 Classical and Time-Varying Autoregression
Time series modeling is a foundational tool in transportation, climate science, econometrics, and other fields [8, 9, 10]. Classical models such as autoregression (AR), moving average (MA), and their combinations (e.g., ARIMA) are widely used for capturing temporal dependencies and seasonal structures in univariate time series data [4, 5, 11]. Among these, AR models remain popular due to their simplicity and interpretability. More recently, applications have demanded models that adapt to non-stationary and time-varying systems. Time-varying AR extends classical AR by allowing coefficients to change over time [12], often incorporating structural or smoothness constraints to ensure interpretability and stability [13]. AR has also been generalized to multivariate and spatial-temporal settings, such as Vector Autoregressive (VAR) models [14, 15]. These ideas have been successfully applied in finance, neuroscience, and dynamic mobility systems.
@@ 3.2 Time Series Autoregression
Fig. 2 illustrates the coefficient vector $\boldsymbol{w}$ estimated for the Chicago ridesharing time series using standard least squares. The resulting vector is dense, with both positive and negative coefficients. A large positive $w_{k}$ at lag $k=168$ reflects the strong weekly periodicity of the time series. However, this dense representation makes it difficult to identify which lags are most important and to quantify periodic structure precisely. This motivates the need for sparse and interpretable AR models.
Figure 2: Estimated AR coefficients $\boldsymbol{w}\in\mathbb{R}^{d}$ for the Chicago ridesharing time series in Fig. 1, with order $d=168$.
@@ 4 Sparse Autoregression
This section introduces Sparse Autoregression (SAR) with $\ell_{0}$-norm induced sparsity constraints for identifying dominant auto-correlations in time series. We first describe the modeling framework and associated statistical guarantees. We then present an MIO formulation for solving the resulting optimization problem exactly and compare solution quality across algorithms.
@@ 4.1 Model Description
Next, we study the statistical learnability of a more general version of Problem (4)—excluding the non-negativity constraint. The following result, which we prove in Appendix A, relies on standard arguments [31].
Theorem 1 (Excess Risk Bound for SAR).
Let $\tilde{\boldsymbol{x}}=\boldsymbol{A}\boldsymbol{w}^{\star}+\boldsymbol{\epsilon}$, where:
  • $\boldsymbol{A}\in\mathbb{R}^{(T-d)\times d}$ is fixed and deterministic,
  • $\boldsymbol{w}^{\star}\in\mathbb{R}^{d}$ satisfies $\|\boldsymbol{w}^{\star}\|_{0}\leq\tau<d/2$,
  • $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\sigma^{2}\boldsymbol{I}_{d})$ is i.i.d. Gaussian noise.
Let $\hat{\boldsymbol{w}}$ be the solution to Problem (4). Then the expected in-sample prediction error satisfies:
$\mathbb{E}\left[\frac{\|\boldsymbol{A}(\hat{\boldsymbol{w}}-\boldsymbol{w}^{\star})\|_{2}^{2}}{T-d}\right]\lesssim\sigma^{2}\cdot\frac{\tau}{T-d}\left(\log\left(\frac{d}{\tau}\right)+1\right).$
Theorem 1 demonstrates that the error scales gracefully with sparsity level $\tau$ and AR order $d$.
@@ 4.3 Empirical Comparison of Solution Quality
The estimation method has a significant impact on the interpretability and fidelity of SAR models. To highlight this, we compare two approaches for solving Problem (4): (i) a greedy non-negative subspace pursuit (NNSP) algorithm [7], and (ii) the exact MIO formulation from Eq. (5). We use the time series from Fig. 1, with AR order $d=168$ (one week of hourly lags) and sparsity level $\tau=2$. The solution returned by NNSP is:
@@
with objective value $f(\boldsymbol{w})=6.25\times 10^{7}$. The MIO solution is both quantitatively superior (lower error) and qualitatively more interpretable: it identifies lag $k=1$ (local auto-correlation) and lag $k=168$ (weekly periodicity), consistent with domain knowledge. The NNSP solution, by contrast, includes a spurious lag at $k=53$ with negligible weight.
When increasing the sparsity level to $\tau=3$, both solvers return:
@@
Fig. LABEL:dominant_supp_plus_coef summarizes how the selected support sets and coefficient magnitudes evolve as a function of $\tau$. These results confirm that high-quality solutions are critical for interpretability and that MIO offers robust, principled support recovery even when greedy methods fail.
5 Time-Varying Sparse Autoregression
In this section, we now extend the SAR formulation to capture non-stationary dynamics by allowing the coefficients to vary across prescribed time segments. This yields the Time-Varying Sparse Autoregression (TV-SAR) model. To solve the associated MIO problem more efficiently, we also introduce a decomposition-based DVP strategy.
@@ 5.1 Model Description
Time-varying AR models allow the AR coefficients to evolve across time, capturing structural shifts in the data-generating process. We partition the time series into $\Gamma\in\mathbb{Z}^{+}$ time segments, where each segment $\gamma\in[\Gamma]$ contains $T_{\gamma}\in\mathbb{Z}^{+}$ time steps. Let $\boldsymbol{x}_{\gamma}=(x_{\gamma,1},x_{\gamma,2},\dots,x_{\gamma,T_{\gamma}})^{\top}\in\mathbb{R}^{T_{\gamma}}$ denote the segment corresponding to time segment $\gamma$. We model each segment with an independent $d$th-order AR process, with potentially different coefficient vectors $\boldsymbol{w}_{\gamma}\in\mathbb{R}^{d}$:
@@ 5.2 Acceleration with Decision Variable Pruning
The computational cost of solving TV-SAR with MIO increases with the AR order $d$, due to the total number of decision variables $(\Gamma+1)d$, which includes $\Gamma d$ real-valued variables for the AR coefficients $\boldsymbol{w}_{\gamma}$ and $d$ binary variables in $\boldsymbol{z}$. When the sparsity level $\tau$ is much smaller than $d$, most of these variables are expected to be zero. We exploit this by introducing a Decision Variable Pruning (DVP) strategy, which leverages subspace pursuit to pre-select a reduced candidate set of variables before solving the MIO. The strategy proceeds in three steps (illustrated in Fig. 4):
@@
6 Spatially and Time-Varying Sparse Autoregression
In this section, we introduce the Spatially and Time-Varying Sparse Autoregression (STV-SAR) model, a generalization of TV-SAR that accounts for both temporal and spatial variations in AR behavior. While TV-SAR captures time-varying dynamics within a single univariate time series, STV-SAR is designed for settings involving large spatiotemporal panels of time series (e.g., satellite-based climate data), where thousands of spatial locations exhibit their own local dynamics. This additional spatial dimension leads to a more expressive—but also more complex—model. In particular, when the number of spatial locations (expressed in terms of their latitudes and longitudes) $M\times N=1$, STV-SAR reduces to TV-SAR. Conversely, STV-SAR enables us to model heterogeneity across locations while leveraging global sparsity and seasonality structures. To solve the associated MIO problem efficiently, we describe a two-stage optimization scheme that separates the learning processes of the global support set and individual coefficient vectors.
@@ 6.1 Model Description
Considering a collection of time series arranged over an $M\times N$ spatial grid, we let $\boldsymbol{X}_{\gamma,t}\in\mathbb{R}^{M\times N}$ denote the spatial matrix at time $t\in[T_{\gamma}]$ in time segment $\gamma\in[\Gamma]$. Each grid cell $(m,n)$ contains a multivariate time series $\{x_{m,n,\gamma,t}\}_{t\in[T_{\gamma}],\gamma\in[\Gamma]}$. Following the same logic as in Section 5, we model each time series with a $d$-order AR process, with coefficient vectors $\boldsymbol{w}_{m,n,\gamma}\in\mathbb{R}^{d}$, and impose a shared global support set across all spatial locations and time segments. Let $\tilde{\boldsymbol{x}}_{m,n,\gamma}$ and $\boldsymbol{A}_{m,n,\gamma}$ be the lagged response vector and design matrix (as in Eq. (3)), respectively, then the MIO problem is formulated as follows,
@@ 7 Experiments
In this section, we conduct extensive experiments to evaluate the proposed SAR models on real-world time series data. The time-varying ridesharing trip time series dataset in NYC allows one to discover daily and weekly periodicities and their changes across different periods from 2019 to 2023. The climate variable time series datasets, including North America climate variables and sea surface temperature, are spatially- and time-varying, enabling the tasks such as the spatial pattern discovery of yearly seasonality and the identification of system evolution related to climate dynamics.
@@ 7.1 Human Mobility Periodicity
7.1.1 Ridesharing Trip Time Series
In this work, we examine the proposed TV-SAR model on the ridesharing trip data in NYC, ranging from February 2019 to December 2023. In terms of human mobility, trips to airports or leaving airports are closely related to the flight schedule, which usually demonstrate strong daily and weekly periodicities. Fig. LABEL:NYC_JFK_airport_heatmap shows the pickup and dropoff trip time series of John F. Kennedy International Airport in NYC. In Figs. LABEL:NYC_JFK_airport_ts_pickup and LABEL:NYC_JFK_airport_ts_dropoff, the trip time series of each week is visualized as a row of the heatmap, and there are around 260 weeks from 2019 to 2023 in total. As can be seen, pickup trips show perturbation and fluctuation due to the influential factors such as flight delay, baggage claim, and travel time to ridesharing pickup zones. The peak time demanding ridesharing services is the nighttime. In contrast, dropoff trips to the departure area of the airport are uniform across different weeks. There are morning and afternoon peak hours with remarkable dropoff trips to the airport. Figs. LABEL:NYC_JFK_airport_ts_pickup_curve and LABEL:NYC_JFK_airport_ts_dropoff_curve show the daily periodicity of aggregated time series of pickup and dropoff trips, in which the dropoff trips are more daily and weekly periodic than the pickup trips.
TABLE II: Objective function $f(\boldsymbol{w})$ in Eq. (4) on the ridesharing pickup trip time series in John F. Kennedy International Airport. The solution algorithms include NNSP, MIO-DVP, and MIO. The unit of objective function values is ×$\times 10^{7}$. Note that the lowest objective function values are emphasized in bold fonts. The last two rows present the average computational times (in seconds) of the algorithm.

### v2

@@ Abstract
Time series autoregression (AR) is a classical tool for modeling auto-correlations and periodic structures in real-world systems. We revisit this model from an interpretable machine learning perspective by introducing sparse autoregression (SAR), where $\ell_{0}$-norm constraints are used to isolate dominant periodicities. We formulate exact mixed-integer optimization (MIO) approaches for both stationary and non-stationary settings and introduce two scalable extensions: a decision variable pruning (DVP) strategy for temporally-varying SAR (TV-SAR), and a two-stage optimization scheme for spatially- and temporally-varying SAR (STV-SAR). These models enable scalable inference on real-world spatiotemporal datasets. We validate our framework on large-scale mobility and climate time series. On NYC ridesharing data, TV-SAR reveals interpretable daily and weekly cycles as well as long-term shifts due to COVID-19. On climate datasets, STV-SAR uncovers the evolving spatial structure of temperature and precipitation seasonality across four decades in North America and detects global sea surface temperature dynamics, including El Niño. Together, our results demonstrate the interpretability, flexibility, and scalability of sparse autoregression for periodicity quantification in complex time series.
@@ 1 Introduction
Many real-world systems exhibit complex temporal patterns, including periodicity, seasonality, and anomalies. Detecting and understanding these patterns is essential for anticipating system behavior, identifying disruptions, and supporting operational decision-making. In dynamic environments such as urban transportation and climate systems, periodicities can shift due to external factors—policy interventions, demand changes, environmental variability, global events, or extreme climate phenomena—making interpretable, adaptive, and data-driven methods indispensable. A central challenge is to automatically identify dominant periodic components from temporally- and spatially-varying systems, track their evolution over a long-term time period, and distinguish true structural changes from random variability.
Urban transportation systems display strong periodicity driven by commuting patterns, business cycles, and travel demand. Figure 1 shows the regularity of ridesharing activity in Chicago, highlighting its weekly mobility patterns. However, such patterns are not fixed—they evolve with infrastructure changes, economic conditions, and disruptive events. For example, the COVID-19 pandemic in 2020 led to a collapse in established mobility periodicities due to lockdowns and shifts to remote work. These disruptions raise key questions: How do periodic structures evolve over time? Can we systematically quantify such changes in a temporally-varying system? Addressing these questions is critical for planning, forecasting, and adaptive resource allocation.
@@
To address these needs, we develop a unified interpretable machine learning framework for identifying and quantifying temporally-varying periodicity and structural shifts in real-world time series. Our models build on classical autoregression (AR) [4, 5] but incorporate sparse structure via $\ell_{0}$-norm induced sparsity constraints to promote interpretability. Inspired by recent advances in sparse regression [6] and time series convolution [7], we formulate interpretable AR models capable of isolating dominant periodic patterns over time and space. To the best of our knowledge, this represents the first application of exact sparse AR to the task of periodicity quantification in real-world complex time series. Overall, this work makes the following contributions:
@@
  • Temporally-Varying SAR (TV-SAR): We extend SAR to characterize non-stationary time series, enforcing consistent support sets across time segments to enhance interpretability. To improve scalability, we introduce a decision variable pruning (DVP) strategy that narrows the MIO search space using fast greedy approximations such as subspace pursuit.
  • Spatially- and Temporally-Varying SAR (STV-SAR): We propose a scalable model for multidimensional time series that vary over both space and time. A two-stage optimization procedure—global support set selection via MIO, followed by local coefficient estimation via quadratic optimization—makes the method tractable for millions of decision variables.
@@ 2 Literature Review
2.1 Classical and Temporally-Varying Autoregression
Time series modeling is a foundational tool in transportation, climate science, econometrics, and other fields [8, 9, 10]. Classical models such as autoregression (AR), moving average (MA), and their combinations (e.g., ARIMA) are widely used for capturing temporal dependencies and seasonal structures in univariate time series data [4, 5, 11]. Among these, AR models remain popular due to their simplicity and interpretability. More recently, applications have demanded models that adapt to non-stationary and temporally-varying systems. Temporally-varying AR extends classical AR by allowing coefficients to change over time [12], often incorporating structural or smoothness constraints to ensure interpretability and stability [13]. AR has also been generalized to multivariate and spatial-temporal settings, such as Vector Autoregressive (VAR) models [14, 15]. These ideas have been successfully applied in finance, neuroscience, and dynamic mobility systems.
@@ 3.2 Time Series Autoregression
Figure 2 illustrates the coefficient vector $\boldsymbol{w}$ estimated for the Chicago ridesharing time series using standard least squares. The resulting vector is dense, with both positive and negative coefficients. A large positive $w_{k}$ at lag $k=168$ reflects the strong weekly periodicity of the time series. However, this dense representation makes it difficult to identify which lags are most important and to quantify periodic structure precisely. This motivates the need for sparse and interpretable AR models.
Figure 2: Estimated AR coefficients $\boldsymbol{w}\in\mathbb{R}^{d}$ for the Chicago ridesharing time series in Figure 1, with order $d=168$.
@@ 4 Sparse Autoregression
This section introduces Sparse Autoregression (SAR) with $\ell_{0}$-norm induced sparsity constraints for identifying dominant auto-correlations in time series. We first describe the modeling framework; then, we present an MIO formulation for solving the resulting optimization problem exactly and compare solution quality across algorithms.
@@ 4.3 Empirical Comparison of Solution Quality
The estimation method has a significant impact on the interpretability and fidelity of SAR models. To highlight this, we compare two approaches for solving Problem (4): (i) a greedy non-negative subspace pursuit (NNSP) algorithm [7], and (ii) the exact MIO formulation from Eq. (5). We use the time series from Figure 1, with AR order $d=168$ (one week of hourly lags) and sparsity level $\tau=2$. The solution returned by NNSP is:
@@
with objective value $f(\boldsymbol{w})=6.25\times 10^{7}$. The MIO solution is both quantitatively superior (lower error) and qualitatively more interpretable: it identifies lag $k=1$ (local auto-correlation) and lag $k=168$ (weekly periodicity), consistent with domain knowledge. The NNSP solution, by contrast, includes a spurious lag at $k=53$ with negligible weight. As can be seen, the choice of solution algorithms affect not only the objective function of optimization problem but the interpretability of sparse coefficient vector as well.
When increasing the sparsity level to $\tau=3$, both solvers return the coefficient vector such that
@@
Figure LABEL:dominant_supp_plus_coef summarizes how the selected support sets and coefficient magnitudes evolve as a function of $\tau$. These results confirm that high-quality solutions are critical for interpretability and that MIO offers robust, principled support recovery even when greedy methods fail.
5 Temporally-Varying Sparse Autoregression
In this section, we now extend the SAR formulation to capture non-stationary dynamics by allowing the coefficients to vary across prescribed time segments. This yields the Temporally-Varying Sparse Autoregression (TV-SAR) model. To solve the associated MIO problem more efficiently, we also introduce a decomposition-based DVP strategy.
@@ 5.1 Model Description
Temporally-varying AR models allow the AR coefficients to evolve across time, capturing structural shifts in the data-generating process. We partition the time series into $\Gamma\in\mathbb{Z}^{+}$ time segments, where each segment $\gamma\in[\Gamma]$ contains $T_{\gamma}\in\mathbb{Z}^{+}$ time steps. Let $\boldsymbol{x}_{\gamma}=(x_{\gamma,1},x_{\gamma,2},\dots,x_{\gamma,T_{\gamma}})^{\top}\in\mathbb{R}^{T_{\gamma}}$ denote the segment corresponding to time segment $\gamma$. We model each segment with an independent $d$th-order AR process, with potentially different coefficient vectors $\boldsymbol{w}_{\gamma}\in\mathbb{R}^{d}$:
@@ 5.2 Acceleration with Decision Variable Pruning
The computational cost of solving TV-SAR with MIO increases with the AR order $d$, due to the total number of decision variables $(\Gamma+1)d$, which includes $\Gamma d$ real-valued variables for the AR coefficients $\boldsymbol{w}_{\gamma}$ and $d$ binary variables in $\boldsymbol{z}$. When the sparsity level $\tau$ is much smaller than $d$, most of these variables are expected to be zero. We exploit this by introducing a Decision Variable Pruning (DVP) strategy, which leverages subspace pursuit to pre-select a reduced candidate set of variables before solving the MIO. The strategy proceeds in three steps (illustrated in Figure 4):
@@
6 Spatially and Temporally-Varying Sparse Autoregression
In this section, we introduce the Spatially and Temporally-Varying Sparse Autoregression (STV-SAR) model, a generalization of TV-SAR that accounts for both temporal and spatial variations in AR behavior. While TV-SAR captures temporally-varying dynamics within a single univariate time series, STV-SAR is designed for settings involving large spatiotemporal panels of time series (e.g., satellite-based climate data), where thousands of spatial locations exhibit their own local dynamics. This additional spatial dimension leads to a more expressive—but also more complex—model. In particular, when the number of spatial locations (expressed in terms of their latitudes and longitudes) $M\times N=1$, STV-SAR reduces to TV-SAR. Conversely, STV-SAR enables us to model heterogeneity across locations while leveraging global sparsity and seasonality structures. To solve the associated MIO problem efficiently, we describe a two-stage optimization scheme that separates the learning processes of the global support set and individual coefficient vectors.
@@ 6.1 Model Description
Considering a collection of time series arranged over an $M\times N$ spatial grid, we let $\boldsymbol{X}_{\gamma,t}\in\mathbb{R}^{M\times N}$ denote the spatial matrix at time $t\in[T_{\gamma}]$ in time segment $\gamma\in[\Gamma]$. Each grid cell $(m,n)$ contains a multivariate time series $\{x_{m,n,\gamma,t}\}_{t\in[T_{\gamma}],\gamma\in[\Gamma]}$, in which any $\gamma$th univariate time series is of length $T_{\gamma}$. Following the same logic as in Section 5, we model each time series with a $d$-order AR process, with coefficient vectors $\boldsymbol{w}_{m,n,\gamma}\in\mathbb{R}^{d}$, and impose a shared global support set across all spatial locations and time segments. Let $\tilde{\boldsymbol{x}}_{m,n,\gamma}$ and $\boldsymbol{A}_{m,n,\gamma}$ be the lagged response vector and design matrix (as in Eq. (3)), respectively, then the MIO problem is formulated as follows,
@@ 7 Experiments
In this section, we evaluate the proposed sparse autoregression (SAR) models through a series of experiments on real-world time series datasets. We consider two distinct domains: urban mobility and climate science. First, we analyze a temporally-varying dataset of ridesharing trips in NYC, which exhibits rich daily and weekly periodicities. This dataset spans multiple years (2019–2023), enabling us to detect how temporal patterns evolve across pre-pandemic, pandemic, and post-pandemic periods. Second, we study large-scale climate datasets, including North American atmospheric variables and global sea surface temperatures. These datasets are both spatially and temporally structured, allowing us to uncover the geographical organization of seasonal effects and trace how climate patterns shift across time.
@@ 7.1 Human Mobility Periodicity
This case study focuses on identifying and tracking periodic patterns in human mobility using ridesharing data in NYC from 2019 to 2023. Due to daily and weekly commuting cycles and disruptions such as COVID-19, the underlying temporal structure is both periodic and nonstationary. We demonstrate how the proposed TV-SAR model captures this evolving structure and compare alternative solution approaches for model estimation.
7.1.1 Dataset and Motivation for TV-SAR
We apply the proposed TV-SAR model to NYC ridesharing data spanning February 2019 to December 2023, focusing on mobility patterns at John F. Kennedy International Airport. Airport-related trips are known to follow strong daily and weekly rhythms due to fixed flight schedules.
Figures LABEL:NYC_JFK_airport_ts_pickup and LABEL:NYC_JFK_airport_ts_dropoff visualize the weekly evolution of pickup and dropoff activities, with each heatmap row corresponding to one week (approximately 260 in total). Pickup trips show higher variability due to factors such as flight delays, baggage collection, and access constraints to pickup zones; demand peaks in the evening hours. In contrast, dropoff trips exhibit clearer regularity, with consistent morning and afternoon peaks aligned with flight departures. Aggregated time series in Figures LABEL:NYC_JFK_airport_ts_pickup_curve and LABEL:NYC_JFK_airport_ts_dropoff_curve further confirm that dropoff trips display stronger daily and weekly periodicity compared to pickups.
Furthermore, as shown in Figure LABEL:NYC_JFK_airport_heatmap, the ridesharing pickup and dropoff time series display long-term temporal variability. Notably, there is a sharp decline in trip volumes in 2020 due to the COVID-19 pandemic, followed by a gradual recovery through the end of 2021. These observations suggest that the periodic structure of the ridesharing time series is not only rich but also subject to external disruptions. This motivates the use of a temporally-varying model such as TV-SAR to flexibly capture changes in autoregressive patterns across months and years.
7.1.2 Model Training: Comparison of Estimation Algorithms
We compare three algorithms for solving Problem (4) on the ridesharing pickup time series over five years: NNSP, MIO-DVP, and MIO. All methods are tested at two sparsity levels, $\tau=4$ and $\tau=6$. For MIO-DVP, we additionally experiment with different pruning thresholds $\tau_{0}=5$ and $\tau_{0}=10$ to examine how the preselected support set size influences solution quality.
Table II reports the objective function values under each setting. As expected, the MIO algorithm consistently achieves the best performance, finding globally optimal solutions. NNSP serves as a baseline and exhibits inferior solution quality, particularly at higher sparsity levels. The performance of MIO-DVP lies in between: with an effective pruning threshold (e.g., $\tau_{0}=10$), it nearly matches the performance of full MIO while incurring significantly lower computational cost.
TABLE II: Objective function $f(\boldsymbol{w})$ in Eq. (4) on the ridesharing pickup trip time series at JFK. The solution algorithms include NNSP, MIO-DVP, and MIO. The unit of objective function values is ×$\times 10^{7}$. Note that the lowest objective function values are emphasized in bold fonts. The last two rows present the average computational times (in seconds).